OpenStudy (anonymous):

John bought 15 apples for Rs.10 and sold them at the rate of 12 apples for Rs.12. What is the percentage of profit made by him

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@UnkleRhaukus ,@Hero

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

\[\frac{\frac{12}{12}-\frac{10}{15}}{\frac{10}{15}}=0.5=50 \% \]